OpenStudy (openstudyrocks5*):

Make a table and graph the function. The Rule is y= x + 3. Click on this question and look at comments to see the answers... Please Help :) Will make best answer for the best answer.

OpenStudy (openstudyrocks5*):

V

OpenStudy (openstudyrocks5*):

@MARC_

OpenStudy (openstudyrocks5*):

@iGreen.

OpenStudy (anonymous):

y = x + 3 The slope here is '1', and that's positive, so the line will rise from left to right..that takes out options A and B. So we're left with C and D.

OpenStudy (anonymous):

Um, C and D are the same.

OpenStudy (anonymous):

Oh wait, nevermind.

OpenStudy (anonymous):

Let's plug in -1 for 'x' into the equation and see what we get. \(y = x + 3\) \(y = -1 + 3\) What's -1 + 3?

OpenStudy (openstudyrocks5*):

-1 + 3 = 2

OpenStudy (openstudyrocks5*):

@iGreen. You there?

OpenStudy (anonymous):

Yes, so when x = -1, y = 2.

OpenStudy (anonymous):

So our answer is C, because D says when x = -1, y = 6.

OpenStudy (anonymous):

ok you need help

OpenStudy (openstudyrocks5*):

Yeah I was thinking it was C. anyway :) Thanks for your help IGreen. I may have more questions.

OpenStudy (anonymous):

\(\bbox [10pt, lime, border:5pt solid black]{\Huge\cal\color{red}\succ\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!\color{red}\prec}}}\)

OpenStudy (anonymous):

ok

OpenStudy (openstudyrocks5*):

Thank you IGreen :). You get the best answer and the only answer lol

OpenStudy (anonymous):

good job

OpenStudy (openstudyrocks5*):

This Question is Now Closed

OpenStudy (anonymous):

yay